SRAM costs about $500 per GB while 16 GB of DRAM costs $100 today. consider scenario where the SRAM has an average access time of 2 ns while the DRAM has an average access time of 50 ns. if we are able to design a cache memory system with 1 MB of SRAM, 16 GB of DRAM and hit rate of 90% what would be the improvement in price/performance ratio?
Answer #1

Answer : price for 1MB SRAM = $ 500/1024

Performance SRAM = 2*0.9

Price 16 GB DRAM = $100

Performance of DRAM = 50*0.9

Improvement in price /performance = 100/(50*0.9) * (1024/500) * 2* 0.9=( 100 * 1024 *2)/ (50*500) = 8.192